### v1

# Anomalous and nonlinear Hall effects in two-dimensional CrPS4

Abstract
Few-layer CrPS4 is a two-dimensional (2D) magnetic material with excellent stability in ambient environment, which attracted significant interest in recent research. Here, via first-principles calculations, we show that 2D CrPS4 hosts a variety of anomalous Hall transport phenomena, owing to its layer-dependent magnetism and symmetry character. Monolayer CrPS4 can display a sizable linear anomalous Hall effect, while its nonlinear anomalous Hall response is forbidden. In contrast, bilayer CrPS4 can produce pronounced intrinsic nonlinear anomalous Hall response from Berry-connection polarizability, in the absence of linear anomalous Hall effect. We clarify that the large peaks in these responses originate from gapped Dirac points in the band structure. Furthermore, we show that linear anomalous Hall effect can be induced and controlled in bilayer CrPS4 by gate electric field or in-plane magnetic field, which break the spacetime inversion symmetry. Our findings unveil the interesting layer-dependent Hall transport physics in 2D CrPS4 magnets, suggesting its potential in electronic and spintronic device applications.
I Introduction
The various Hall effects are of fundamental importance in condensed matter physics. A prominent example is the (linear) anomalous Hall effect (AHE), where a Hall voltage is induced by a longitudinal current flow in a magnetic material [1]. An important advance in the past two decades is the recognition that this effect contains an intrinsic contribution, i.e., an response determined solely by the material’s band structure, which manifests an intriguing band geometric property, the Berry curvature of electronic bands [2, 3]. AHE is a time-reversal-odd ($\mathcal{T}$-odd) effect, meaning that reversal of magnetic moments in the system (corresponding to a time-reversal operation) must flip the sign of the response coefficient. The $\mathcal{T}$-odd Hall effect was also discovered in nonlinear response of magnets, giving a Hall current quadratic in the driving $E$ field, i.e., $j_{\text{H}}\propto E^{2}$ [4, 5, 6]. Importantly, this nonlinear AHE also contains an intrinsic contribution, which is determined by another band geometric quantity, the Berry-connection polarizability (BCP) [6, 7], and can be further connected to the quantum metric of band structure [5, 8, 9]. Note that this effect is distinct from the $\mathcal{T}$-even nonlinear AHE that has been widely studied in nonmagnetic materials and does not have an intrinsic contribution [10, 11, 12, 13]. Recently, such intrinsic nonlinear AHE has been experimentally detected in a few magnets, such as few-layer MnBi2Te4 [14, 8] and Mn3Sn [15].
Meanwhile, a focus of current materials research is to explore two-dimensional (2D) magnets [16, 17, 18]. Several material classes, like CrI3 [19], Fe3GeTe2 [20, 21], Cr2Ge2Te6 [22], and MnBi2Te4 [23, 24], have been realized. Due to the reduced dimensionality, a big advantage of 2D magnets is that their properties can be flexibly tuned, e.g., by gating, doping, strain, external fields, and hetero-structure engineering. Nevertheless, it was noted that most of 2D magnets studied so far are not stable in ambient environment, posing difficulties for applications. In this regard, the recently realized 2D magnet CrPS4 starts to attract increasing attention because of its excellent air stability [25]. It was shown that its magnetic ordering can persist down to monolayer limit [25]. The physical properties of CrPS4, such as spin dynamics, layer-dependent magnetism, optical response, electric and thermal transport properties, have been studied in several recent works [26, 27, 28, 29, 30, 25, 31, 32, 33, 34, 35, 36, 37]. As a semiconductor, the type and density of its charge carriers can be conveniently controlled by gating [35, 36]. Devices based on 2D CrPS4, such as field effect transistors and magnetic tunnel junctions, have been studied in experiment [36, 38, 37].
Here, we note that 2D CrPS4 actually offers a good platform to study a variety of Hall effects. First, 2D CrPS4 with odd number of layers is ferromagnetic (FM), which allows linear AHE. Second, even-layer CrPS4 has a compensated antiferromagnetic (AFM) ordering, which preserves spacetime inversion ($\mathcal{PT}$) symmetry. This forbids the linear AHE. However, the intrinsic nonlinear AHE is permitted. It should be noted that the symmetry condition of this nonlinear effect in 2D systems is a bit stringent: It requires the absence of any out-of-plane rotation axis [5, 6]. CrPS4 naturally satisfies this condition due to its monoclinic crystal structure. Third, although linear AHE is forbidden in the ground state of even-layer CrPS4, external fields may break the $\mathcal{PT}$ symmetry and generate a field-induced AHE.
Motivated by these considerations, in this work, we perform a study of the various Hall responses in 2D CrPS4 via first-principles calculations. We focus on monolayer and bilayer CrPS4 as representatives. For monolayer CrPS4, we show that it has a sizable anomalous Hall conductivity in the case of hole doping. Its value can be strongly influenced by the magnetization direction. For bilayer CrPS4, we show that the intrinsic nonlinear AHE is determined by a single response coefficient, which can achieve a large value $\sim 1$ mA/V2 under hole doping. The origin of the large contribution in BCP is traced to a band anti-crossing region in valence bands. Furthermore, to induce linear AHE in bilayer CrPS4, we break the $\mathcal{PT}$ symmetry in two ways. The first is by a gate electric field (in the out-of-plane direction). The resulted net AHE response has a sign determined by the gate field direction, and the Hall current has a dominant distribution in one of the layers. The second way is to apply an in-plane magnetic field, which causes canting of the local moments and leads to an in-plane magnetization. We show that the resulting AHE can be quite large under moderate applied field strength. The two methods are also discussed from the perspective of layer Hall effect. Our results reveal interesting Hall transport physics in the 2D magnet CrPS4, which may find potential applications in novel electronic and spintronic devices.
II Computation method
[…]
Figure 1: (a) Top view of a CrPS4 layer. The black rectangle marks the conventional cell. (b) Lattice structure of bulk CrPS4. It has a $A$-type AFM ground state. The local moments are in the out-of-plane direction, as indicated by the red arrows in the figure.
III Structure and magnetic ordering
Bulk CrPS4 is a vdW layered material with monoclinic crystal structure belonging to the $C2/m$ space group (see Fig. 1). It was first synthesized and studied in the 1970s [48]. A recent experimental characterization reported that its lattice parameters (for the conventional cell) are $a=10.856\,\text{\AA}$, $b=7.247\,\text{\AA}$, and $c=6.135\,\text{\AA}$ [29]. As shown in Fig. 1(a), in a layer, each Cr cation is surrounded by six S anions, forming a CrS6 octahedron with slight Jahn-Teller distortion. These CrS6 octahedra form quasi-1D chains along $y$ direction in Fig. 1(a), which are further inter-connected by PS4 tetrahedra, resulting in a rectangle lattice in the 2D plane. Due to the relatively weak vdW interaction, a slight slip along a axis direction occurs between neighboring layers, causing angle $\beta=88.116^{\circ}$, which deviates slightly from $90^{\circ}$ (see Fig. 1(b)). The bulk structure of CrPS4 preserves the spatial inversion symmetry $\mathcal{P}$. However, due to the in-plane anisotropy as described above, there is no vertical rotation axis (only a twofold axis along $y$ exists).
The magnetic ground state of CrPS4 is well established in previous studies [25, 29]. The magnetism comes mainly from the Cr3+ ions with a magnetic moment about $3.0\mu_{B}$. The magnetic coupling within each layer is FM and the moments are in the out-of-plane direction; whereas the coupling between neighboring layers is AFM. This results in an overall $A$-type AFM ordering for bulk CrPS4, with Néel vector along $z$ axis, as shown in Fig. 1(b). The Néel temperature reported in experiment is about 35 to 38 K [28, 30, 29, 25, 38].
As a vdW layered material, 2D few layers can be easily exfoliated from bulk CrPS4. Experimentally, 2D CrPS4 with thickness down to monolayer has been successfully fabricated, via mechanical exfoliation or thermal atomic layer etching method [25]. Compared to other 2D magnets, a big advantage of this material is its good stability in air. It was shown that the magnetism is robust down to monolayer limit, and the ground state exhibits a layer-dependent alternation between two configurations. The odd layers are uncompensated antiferromagnet with a net magnetization along $z$ (FM for monolayer), whereas the even layers feature a fully compensated AFM state. The magnetic transition temperature shows a slight decrease when thickness decreases. The transition temperature for monolayer is found to be $\sim 22$ K [25]. Taking into account the ground state magnetic configuration, the odd layers have $2^{\prime}/m^{\prime}$ magnetic point group, whereas the even layers belong to $2/m^{\prime}$. Their key difference is that the odd layers preserve $\mathcal{P}$ (not $\mathcal{PT}$), but the even layers preserve $\mathcal{PT}$ (not $\mathcal{P}$). This has important consequences on the type of Hall effects allowed in the system, as we shall discuss in a while.
Figure 2: (a) Brillouin zone and (b) calculated band structure for bulk CrPS4.
The band structure for bulk CrPS4 in the AFM ground state obtained from our DFT calculation is plotted in Fig. 2(b). The result shows that CrPS4 is a magnetic semiconductor. The obtained bandgap value $\sim$1.21 eV agrees very well with the recent experimental result ($\sim$1.31 eV) [28]. Our calculated magnetic moment $\sim 3.0\mu_{B}$ on Cr site also agrees with several previous experiments [25, 29]. These support the validity of our computation method. In the case of few layers, we find that they are also magnetic semiconductors. The local magnetic moments remain nearly unchanged compared to the bulk. These features are consistent with previous studies. In the following, we shall focus on monolayer and bilayer CrPS4 to study their various Hall responses.
IV Anomalous Hall response in monolayer
Due to $A$-type magnetic ordering, odd-layer CrPS4 has an uncompensated net magnetization. For monolayer CrPS4, this reduces to the FM ordering, with all magnetic moments in the out-of-plane direction ($z$ direction in our setup). The band structure for this ground state is plotted in Fig. 3(c) The result shows a semiconductor with a bandgap $\sim$1.20 eV, which agrees with previous studies [25].
As we noted above, the magnetic point group for monolayer CrPS4 (with FM ordering) is $2^{\prime}/m^{\prime}$, which consists of two elements: $\mathcal{P}$ and $\mathcal{T}\mathcal{M}_{y}$ ($\mathcal{M}_{y}$ is the mirror normal to $y$). This symmetry allows for a nonzero $\sigma_{xy}$ for linear AHE. On the other hand, any nonlinear charge transport, including nonlinear Hall effect, is forbidden.
Figure 3: (a) Lattice structure of monolayer CrPS4, which has FM ordering with out-of-plane magnetization. (b) Brillouin zone of the monolayer. (c) Band structure for monolayer CrPS4 in the FM ground state.
Here, we compute the intrinsic AHE response. The corresponding anomalous Hall conductivity is given by [49, 50]
$\sigma_{xy}=\frac{e^{2}}{\hbar}\int[d\bm{k}]f_{n\bm{k}}\Omega_{z}^{n}(\bm{k}),$ (1)
where $[d\bm{k}]$ is the short notation of $\sum_{n}d\bm{k}/(2\pi)^{2}$, $n$ is the band index, $f_{n\bm{k}}$ is the Fermi-Dirac distribution function, and
$\Omega_{z}^{n}(\bm{k})=-2\hbar^{2}\text{Im}\sum_{m\neq n}\frac{(v_{x})_{nm}(v_{y})_{mn}}{(\varepsilon_{n}-\varepsilon_{m})^{2}}$ (2)
is the Berry curvature of Bloch band structure, with $\varepsilon_{n}$ being the band energy and $(v_{i})_{mn}$ being the interband velocity matrix element. For simple notations, we suppress the $k$ dependence in Eq. (2) above. From Eq. (1), one can see that this AHE response is intimately connected to Berry curvature of occupied states.
Figure 4: (a) Calculated intrinsic anomalous Hall conductivity for monolayer CrPS4 as a function of chemical potential. (b) Distribution of Berry curvature for states below $\mu=-0.32$ eV. (c) The hot spots in (b) correspond to the gapped Dirac points in band structure. The red arrow indicates one point. The other one is related to this one by inversion symmetry. (d) The result of intrinsic anomalous Hall conductivity when the magnetization is rotated to be along the $x$ direction.
In Fig. 4(a), we show the calculation result of $\sigma_{xy}$ as a function of the chemical potential $\mu$. One can observe several peaks in both hole-doped and electron-doped regions. The peak at $\mu\sim-0.32$ eV can reach up to $\sim-600$ S/cm, which is a quite large value. To understand the origin of this peak, we calculate the Berry curvature distribution in $k$ space summed over states below $-0.32$ eV. The result is shown in Fig. 4(b). One can see that large Berry curvature comes from two hot spots sitting on the $\Gamma$-$K$ path. These hot spots correspond to the small-gap regions indicated in Fig. 4(c). It should be noted that here, the two bands do not cross, instead, they form an anti-crossing with a small gap $\sim 0.014$ eV. Such anti-crossing may be considered as a 2D gapped Dirac point. From symmetry analysis, we construct the following $k\cdot p$ effective model for this gapped Dirac point:
$H_{\text{eff}}=wk_{y}+v_{x}k_{x}\tau_{y}+v_{y}k_{y}\tau_{z}+m\tau_{x},$ (3)
where momentum $k$ is measured from the Dirac point, $\tau$’s are Pauli matrices, $w$, $v_{i}$, and $m$ are real model parameters. It is known that Berry curvature is concentrated around such points, where the interband coherence is strong. This explains the feature that we observe in Fig. 4(c).
The magnetic anisotropy in 2D CrPS4 is not very large. It was shown in experiment that the magnetic moment direction can be rotated by relatively small applied magnetic field $\sim 90$ mT [25]. Hence, we also consider the AHE response in monolayer when the magnetization is rotated to an in-plane direction. In Fig. 4(d), we show the calculated $\sigma_{xy}$ for magnetization along the $x$ direction. Again, several peaks can be observed in the figure. The peak values are comparable to those in Fig. 4(a). The first peak in valence band has a value $\sim 370$ S/cm. Furthermore, for magnetization along $y$ direction, $\sigma_{xy}$ is identically zero. This is because in this configuration, the vertical mirror $\mathcal{M}_{y}$ symmetry is preserved, which forbids the AHE.
V nonlinear Hall response in bilayer
For the even-layer CrPS4, the state above Néel temperature $T_{N}$ possesses both $\mathcal{P}$ and $\mathcal{T}$ symmetries, which prohibit linear AHE and any second order current response. Below $T_{N}$, the AFM state preserves $\mathcal{PT}$ symmetry, so that the linear AHE is still suppressed, however, the nonlinear Hall response starts to appear. This nonlinear response can be expressed as
$j_{a}=\chi_{abc}E_{b}E_{c},$ (4)
where $\chi_{abc}$ is the nonlinear conductivity tensor, the subscripts label the Cartesian components, and the Einstein summation convention is assumed here. Here, we focus on the intrinsic contribution to nonlinear Hall transport, which is determined solely by the band structure property. It should be noted that the extrinsic contribution from Berry curvature dipole is forbidden by the $\mathcal{PT}$ symmetry here [10].
Figure 5: (a) Lattice structure of bilayer CrPS4, which has $A$-type AFM ground state. (b) Calculated band structure for bilayer CrPS4 in AFM ground state.
The theory of intrinsic nonlinear Hall effect has been developed in Ref [5, 4, 6]. The corresponding nonlinear conductivity tensor can be expressed as
$\chi_{abc}=\frac{e^{2}}{\hbar}\int[d\bm{k}]f_{n\bm{k}}\Lambda^{n}_{abc}(\bm{k})$ (5)
where
$\Lambda^{n}_{abc}(\bm{k})=\partial_{b}G_{ac}^{n}-\partial_{a}G_{bc}^{n}$ (6)
is the $k$-resolved BCP dipole, and
$G_{ab}^{n}(\bm{k})=2e\hbar^{2}\mathrm{Re}\sum_{m\neq n}\frac{(v_{a})_{nm}(v_{b})_{mn}}{(\varepsilon_{n}-\varepsilon_{m})^{3}}$ (7)
is the BCP tensor, and $\partial_{a}\equiv\partial_{k_{a}}$. Like Berry curvature, BCP is a gauge-invariant quantity. Physically, it characterizes the shift of a electron wavepacket enter by an applied electric field [4]. Near hot spots formed by a pair of bands, BCP can be further connected to the quantum metric of band structure [9].
In the following, we focus on the bilayer CrPS4 as a representative. The calculated band structure is plotted in Fig. 5(b). It should be noted here that each band is doubly spin degenerate due to the $\mathcal{PT}$ symmetry. As for the nonlinear conductivity tensor, we note that the $2/m^{\prime}$ symmetry of the system dictates that the intrinsic response is characterized by a single independent component, with
$\chi_{yxx}=-\chi_{xyx}.$ (8)
For an in-plane $E$ field in the direction $\bm{E}=E(\hat{x}\cos{\phi}+\hat{y}\sin{\phi})$, where $\phi$ is the polar angle measured from $x$ axis, the resulting Hall current $j_{\text{H}}$ in the direction normal to $E$ field can be expressed as
$j_{\text{H}}=\chi_{\text{H}}(\phi)E^{2},\qquad\chi_{\text{H}}(\phi)=\chi_{yxx}\cos\phi.$ (9)
Figure 6: (a) Calculated intrinsic nonlinear Hall conductivity $\chi_{yxx}$ as a function of chemical potential for bilayer AFM CrPS4. (b) $k$-space distribution of BCP dipole for states below $\mu=-0.3$ eV. The hot spots on $\Gamma$-$K$ path correspond to the two gapped Dirac points marked in the enlarged band structure in (c).
In nonlinear Hall experiment, one typically fabricate a disk shaped device and multiple leads, such that the driving current can be applied along various in-plane directions while measuring the Hall response in the transverse direction [12, 51]. Equation (9) tells us the angular dependence of the measured Hall signal should exhibit a simple cosine dependence, with a $2\pi$ periodicity.
Next, we compute the nonlinear conductivity component $\chi_{yxx}$ from our DFT result. The result is shown in Fig. 6(a). One observes that the value on the electron-doping side can reach $\sim 0.4$ mA/V2. For hole doping, the peak at $\mu=-0.3$ eV is even larger, reaching $\sim 2$ mA/V2, which is quite large. It is comparable to the one measured in few-layer MnBi2Te4 [8, 52].
To reveal the origin of this peak, we plot the distribution of $k$-resolved BCP dipole $\Lambda_{yxx}(\bm{k})$ in BZ by summing over all occupied states up to $\mu=-0.3$ eV. The result is plotted in Fig. 6(b). One can see a pronounced dipole like distribution in a small region around a point on $\Gamma$-$K$ ($k_{y}>0$) path. This corresponds to a band anti-crossing point as indicated in Fig. 6(c). By analyzing the band structure around this point, we identify it as a gapped Dirac point. However, different from that in Fig. 4(d), the bands are doubly degenerate. The effective model for this point from symmetry analysis takes the following form:
$\begin{split}H_{\text{eff}}=wk_{y}+v_{x}k_{x}(\sigma_{x}\cos\alpha+\sigma_{y}\sin\alpha)\tau_{y}\\ +v_{y}k_{y}\sigma_{z}\tau_{y}+m\tau_{x},\end{split}$ (10)
where the wavevector $k$ is measured from the point, $\sigma$ and $\tau$ are two sets of Pauli matrices, $w$, $v$’s, $\alpha$, and $m$ are real model parameters. The last term in (10) opens a small gap $\sim 0.007$ eV between the two doubly degenerate bands. As shown in Ref. [6, 5], such gapped Dirac point tends to enhance BCP hence makes significant contribution to the nonlinear Hall response. In Fig. 6(c), there is another hot spot on the $k_{y}<0$ portion of $\Gamma$-$K$ path, with a comparatively weaker BCP dipole distribution. This corresponds to another gapped Dirac point at energy slightly below $-0.3$ eV (see Fig. 6(b)). The form of its effective model is the same as Eq. (10), since the local symmetry is the same.
VI Layer Hall effect in bilayer
Figure 7: (a) Calculated layer Hall conductivity in bilayer CrPS4. Here, we focus on the peak structure in valence band. (b) Schematic figure showing the application of a gate electric field to bilayer CrPS4. This breaks the $\mathcal{PT}$ symmetry and generates a finite anomalous Hall response. (c) Layer-resolved Hall conductivities under a gate field of $0.02$ V/nm. (d) The resulting anomalous Hall conductivity versus gate field at a fixed chemical potential of $-0.25$ eV.
We have shown that a large AHE can be realized in monolayer CrPS4. The bilayer is made of two monolayers with opposite magnetization. Hence, the vanishing of AHE in bilayer can be intuitively understood as the exact cancellation of the AHE responses of the two monolayers. From this perspective, one may consider the bilayer system actually hosts a layer Hall effect [53], i.e., the AHE response is finite in each layer channel and is opposite between the two channels. This is analogous to concepts of the spin Hall effect and valley Hall effect, with layer index being the binary degree of freedom.
Similar to spin or valley Berry curvature, we may define a layer Berry curvature [54] for bilayer CrPS4:
$\Omega_{z}^{n,\ell}(\bm{k})=-2\hbar^{2}\text{Im}\sum_{m\neq n}\frac{(v_{x}^{\ell})_{nm}(v_{y})_{mn}}{(\varepsilon_{n}-\varepsilon_{m})^{2}},$ (11)
where $\ell=T,B$ is the binary layer index for top and bottom layers, $(v_{x}^{\ell})_{nm}$ is the interband matrix element for the operator $\hat{v}_{x}^{\ell}=\{\hat{P}_{\ell},\hat{v}_{x}\}/2$, i.e., the velocity operator combined with the projector $\hat{P}_{\ell}$ into layer $\ell$. The response of AHE in each layer is characterized by
$\sigma_{xy}^{\ell}=\frac{e^{2}}{\hbar}\int[d\bm{k}]f_{n\bm{k}}\Omega_{z}^{n,\ell}(\bm{k}).$ (12)
One can easily see that the total anomalous Hall conductivity
$\sigma_{xy}=\sigma_{xy}^{T}+\sigma_{xy}^{B},$ (13)
must vanish. However, the layer Hall conductivity
$\sigma_{xy}^{L}=\sigma_{xy}^{T}-\sigma_{xy}^{B}$ (14)
can be nonzero.
In Fig. 7(a), we show the calculated layer Hall conductivity for bilayer CrPS4. One can see its peak value can reach $\sim 220$ S/cm on the hole doping side. This indicates that although the net charge current flow vanishes, the AHE response within each layer is still quite large. Therefore, if one can break the symmetry that connects the two layers (here it is $\mathcal{PT}$), it will be possible to generate an imbalance between $\sigma_{xy}^{T}$ and $\sigma_{xy}^{B}$ and induce a finite AHE. A simplest method is by using the gate electric field.
Figure 7(c) shows the calculated layered-resolved and net AHE conductivities for bilayer CrPS4 under a gate field of $0.02$ V/nm. One observes that the gate field shifts the layered-resolved conductivities $\sigma_{xy}^{\ell}$ in different ways, leading to a finite AHE. In Fig. 7(d), we plot the net AHE conductivity as a function of gate field at a fix chemical potential $\mu=-0.25$ eV. One observes the linear scaling at weak field region, and the sign of AHE response can be controlled by the direction of gate field, which is a nice feature for device applications.
VII In-plane anomalous Hall effect in bilayer
Another way to break the $\mathcal{PT}$ symmetry of bilayer CrPS4 is to apply an in-plane magnetic field [55]. (Out-of-plane $B$ field also works, but it will bring in the ordinary Hall effect.) The main effect of this in-plane $B$ field is to tilt the local magnet moments and cause a net in-plane magnetization $\delta M$ along the $B$ field direction. It was found from experiment that the magnetic anisotropy in CrPS4 is relatively weak, so the spin canting can be readily achieved by moderate $B$ field strength [29].
The direction of in-plane $B$ field affects the induced AHE. For field along $y$ direction, the resulting magnetization $\delta M_{y}$ cannot efficiently generate AHE, because it preserves the $\mathcal{M}_{y}$ symmetry for each layer. In comparison, the component $\delta M_{x}$ along $x$ direction is more effective in producing the AHE response. In the following, we shall focus on the this case, with $B$ field causing spin canting in the $x$ direction.
In DFT calculation, we model the spin-canted state by rotating the local moments towards the $x$ direction by a canting angle $\theta$. At a small angle $\theta=30^{\circ}$, the calculated layered-resolved and net AHE conductivities are plotted in Fig. 8(c). One observes that the contributions from the two layers no longer cancel out. A sizable AHE conductivity with peak value $\sim 132$ S/cm can be obtained. In Fig. 8(b), we plot the induced AHE conductivity as a function of angle canting $\theta$ for a fixed $\mu=-0.31$ eV. The curve shows a linear increase at small angles and tends to saturate at large angle. It reaches $\sim 201$ S/cm at angle $\theta=60^{\circ}$.
The magnitude of $B$ field needed to achieve the canting angle $\theta$ can be estimated as $B\approx 2J_{c}S\sin(\theta)/(g\mu_{\mathrm{B}})$. Here, we take the interlayer exchange coupling $J_{c}=0.15$ meV as obtained from experiment [25], and $S=3/2$. In this estimation, the magnetic anisotropy energy is neglected, since it is two order of magnitude smaller than the exchange energy [25]. We find that a canting angle of $5^{\circ}$ can be achieved by a moderate $B$ field of $340$ mT, showing that using in-plane $B$ field is an effective way to control the AHE response in bilayer CrPS4.
Figure 8: (a) Schematic figure showing the application of an in-plane magnetic field to bilayer CrPS4. The field breaks $\mathcal{PT}$ symmetry, tilts the local moments, and induces a finite linear anomalous Hall response. (b) Calculated anomalous Hall conductivity versus spin-canting angle $\theta$ at a fixed chemical potential of $-0.31$ eV. (c) Calculated layer-resolved and total anomalous Hall conductivities for the bilayer when the spin-canting angle $\theta=30^{\circ}$.
VIII Discussion and Conclusion
In this study, we have investigated the rich anomalous Hall transport phenomena in 2D CrPS4 systems. Monolayer and bilayer CrPS4 are used as representatives for demonstration. The general features apply also to other few-layer CrPS4 systems. The odd layers share the same symmetry as monolayer, while the even layers share the same symmetry as bilayer. Hence, linear AHE is expected in odd layers, whereas nonlinear Hall effect is expected in even layers. The linear AHE induced by gate field or in-plane magnetic field should occur in even-layer CrPS4 as well.
Experimentally, 2D CrPS4 has been successfully fabricated via various techniques down to monolayer limit [25, 28]. Field effect transistor devices based on 2D CrPS4 have also been demonstrated [37, 31, 36, 35]. The phenomena predicted here should be detectable via standard transport measurement with a dual gate device setup. For the nonlinear Hall effect, the typical method is to modulate the driving current with a low frequency (less than 100 Hz) and measure the response signal at second harmonic frequency by using the lock-in technique [11, 12]. The angular dependence of the signal (as manifested in Eq. (9)) can be probed by fabricating a disk shaped device with multiple leads to allow driving current (and response signal) exerted (detected) along different in-plane directions.
For odd-layer CrPS4, the nonlinear transport signal should vanish both above and below magnetic transition temperature $T_{c}$. For even layers, the nonlinear signal is forbidden above $T_{c}$ (since the crystal structure preserves $\mathcal{P}$) and appears only when temperature is lowered below $T_{c}$. This could be used as a signature for detecting $T_{c}$ or the parity of number of layers. In addition, we note that for temperature above $T_{c}$, both linear and second-order Hall effects are forbidden by the $\mathcal{T}$ and $\mathcal{P}$ symmetries. The leading order anomalous Hall response should be from the third order [51, 56], which was recently probed in several nonmagnetic materials.
In this work, we focus on the intrinsic responses, which represent inherent material properties. There also exist extrinsic contributions arising from disorder scattering, which vary from sample to sample [57]. Experimentally, a typical method to distinguish the different contributions is to perform a scaling analysis by varying the various disorder strengths. The intrinsic response correspond to a component that is independent of scattering. And first-principles result serves as an important benchmark for such analysis.
In conclusion, we have investigated various anomalous Hall transport phenomena in 2D CrPS4. Due to their low crystal symmetry and layer-dependent magnetism, we find interesting linear and nonlinear AHEs, with sizable conductivities predicted by our first-principles calculations. These effects allow us to probe the intrinsic band geometric quantities, such as Berry curvature, BCP, and quantum metric, of the material system. The demonstrated controllability of Hall transport in 2D CrPS4 also offers possible mechanisms for device operations. Our predications here can be readily tested in experiment, and they may find promising applications in next-generation electronics and spintronics based on 2D magnetic materials.
Acknowledgements.
The authors thank D. L. Deng for valuable discussions. This work is supported by UM Start-up Grant (SRG2023-00033-IAPME, SRG2022-00030-IAPME, and SRG2023-00057-IAPME), UM Multi-Year Research Grant (MYRG-GRG2023-00206-IAPME-UMDF), Science and Technology Development Fund of Macau SAR (0048/2023/RIB2 and 0066/2024/RIA1), National Key R&\&&D Program of China (Grant No. 2022YFA1402600), and NSFC (Grants No. 12174018).